# Confronting Dark Matter Capture Rate with Continuous Gravitational Wave Probe of Local Neutron Stars

### v2

@@ VI Discussion and Conclusion
Studying DM capture by NSs could offer valuable insights into the nature of this elusive substance, shedding light on its properties and interactions with ordinary matter. This research holds promise for advancing our understanding of fundamental physics and unlocking the mysteries of the cosmos. Our motivation for this work is to target the nearby local NSs and study the DM capture rate inside them. Generally, NSs are considered to be very efficient in capturing DM with feeble interactions because of their small radius and strong lower gravitational focus power. But, to this date, only a handful number of NSs have been detected (even though they are very abundant in the galaxy $\approx~{}10^{8}$) and most of them are not very close to Earth (as compared to BDs). This eventually also weakens the bounds on scattering cross-section ($\sigma_{\chi n}$) when our goal is to detect the gamma rays and neutrinos resulting from DM annihilation with space or ground-based detectors. Recently, gravitational wave searches have gained a lot of attention which embarks our interest in the detection of local NSs being the natural source of continuous gravitational wave emission. The search of CGW indicates a substantially rich NS population adjacent to Earth and strengthens the bound on $\sigma_{\chi n}$. For the purpose of this paper, we combine the sensitivity from current and future generation gamma-ray and neutrino detectors which enables us to construct a comprehensive picture of DM signature from local NSs.
@@
  • We interpret the differential flux sensitivity (Fig.1) of our selected gamma-ray and neutrino detectors (see sec.V.1) and establish stacked constraints on $\sigma_{\chi n}$ concerning $m_{\chi}$ by combining the flux limits from 8 nearby NSs (Fig.3) assuming an NFW density profile. Fig.3 indicates that constraints are only obtainable for DM masses $<$ 1 GeV with 8 local NSs, and meaningful bounds can be provided by IceCube-Gen2, IceCube, and TRIDENT. No bounds are obtained for gamma-ray detectors.
  • With the inclusion of a few nearby NSs (0.6 kpc), we determine that no limits can be obtained for DM masses above 1 GeV (Fig. 3). We further explore the potential impact of the expected population of local NSs from CGW searches on our current limits (Fig. 4). We follow a recent study by Reed et al. Reed et al. (2021) which extensively surveyed CGWs originating from unidentified Galactic neutron stars near Earth to establish the number of NSs being surveyed by a given CGW investigation.
@@
  • We derive the bounds on $\sigma_{\chi n}$ from the cumulative emission of the local NS population within 0.6 kpc from Earth (Fig. 5). The cumulative emission expected from approximately $6\times 10^{5}$ NSs indicates that the most stringent bounds would originate from TRIDENT (for neutrinos) for DM masses $m_{\chi}<$ 1000 GeV and from SWGO (for gamma-rays) for DM masses $m_{\chi}>$ 1000 GeV.
  • Figs. 3 and 5, where we compare our obtained limits with experimental bounds from direct detection and astrophysical limits, further demonstrate that our results, even with only 8 nearby NSs, are competitive with existing direct detection limits and some astrophysical bounds.
@@
The study reveals that present and upcoming neutrino and gamma-ray observatories are not very sensitive to DM-nucleon scattering when a few known local NSs within 0.6 kpc in the neighborhood of Earth are considered and can only provide a weak bound for DM mass $\sim 1$ GeV. Improved limits on DM-nucleon scattering cross-section can be achieved with the consideration of about $10^{5}$ local NS population anticipated from the CGW search for a range between low ($\sim$ sub-GeV) to heavy ($\sim$ PeV) DM mass. The corresponding limit on $\sigma_{\chi n}$ with the local NS population is found to be on par with those obtained from DM direct search and indirect search bounds from other astrophysical sources like the Sun or Galactic NS population. Future CGW searches with better sensitivity are expected to further limit the local NS population, which can further constrain the DM-nucleon scattering with better precision.
@@ Acknowledgements.
A.D.B. acknowledges financial support from DST, India, under grant number IFA20-PH250 (INSPIRE Faculty Award). Some part of P.B.’s work was supported by the EOSC Future project which was co-funded by the European Union Horizon Programme call INFRAEOSC-03-2020, Grant Agreement 101017536. The authors would like to thank Dr. Tarak Maity and Dr. Nicholas L. Rodd for their valuable suggestions.

### v3

@@ VI Conclusion
Probing NSs in light of DM capture offers valuable insights into the nature of this elusive substance, shedding light on its properties and interactions with ordinary matter. This research holds promise for advancing our understanding of fundamental physics and unlocking the mysteries of the cosmos. Our motivation for this work is to target the nearby local NSs and study the DM capture rate inside them. Generally, NSs are considered to be very efficient in capturing DM with feeble interactions because of their small radius and strong lower gravitational focus power. But, to this date, only a handful number of NSs have been detected (even though they are expected to be very abundant in the galaxy $\approx~{}10^{8}$) and most of them are not very close to Earth (as compared to BDs). This eventually also weakens the bounds on scattering cross section ($\sigma_{\chi n}$) when our goal is to detect the gamma rays and neutrinos resulting from DM annihilation with space or ground-based detectors. Recently, gravitational wave searches have gained a lot of attention, fueling our interest in the detection of local NSs, being the natural source of CGW emission. The search for CGW indicates a substantially rich NS population adjacent to Earth and strengthens the bound on $\sigma_{\chi n}$. For the purpose of this paper, we combine the sensitivity from current and future generation gamma-ray and neutrino detectors which enables us to construct a comprehensive picture of DM signature from local NSs.
@@
  • We interpret the stacked constraints on $\sigma_{\chi n}$ concerning $m_{\chi}$ by combining the flux limits from 8 nearby NSs (Fig. 3) assuming an NFW density profile. Fig. 3 indicates that constraints are only obtainable for DM masses $<$ 1 GeV with 8 local NSs, and meaningful bounds can be provided by IceCube Gen 2, and TRIDENT.
  • We further explore the potential impact of the expected population of local NSs from CGW searches on our current limits (Fig. 4). With the unknown NS population taken into account from CGW searches by Reed et al.Reed et al. (2021), we conclude that present gamma-ray and neutrino detectors are sensitive to GeV scale DM with $\sigma_{\chi n}\simeq 10^{-45}$ cm2.
@@
  • We derive the bounds on $\sigma_{\chi n}$ from the cumulative emission of the local NS population within 0.6 kpc from Earth (Fig. 5). The cumulative emission expected from approximately $10^{5}$ NSs indicates that the most stringent bounds of $\sigma_{\chi n}~{}\sim~{}10^{-45}$ cm2 originates from TRIDENT (for DM annihilation into neutrinos) for DM masses $m_{\chi}<$ 1000 GeV and from SWGO (for DM annihilation into gamma-rays) for DM masses $m_{\chi}>$ 1000 GeV. The bounds obtained on $\sigma_{\chi_{n}}$ are found to be stronger than the IceCube limits from the Sun and a few orders of magnitude weaker than the H.E.S.S. limit obtained from the Galactic NS population.
VII Discussion
This study upholds a promising avenue for local NS from a data analysis perspective with some limitations that can impact our results. We also want to add that when searching for a DM signal from a local neutron star, the impact of backgrounds for neutrinos and gamma rays is less significant compared to diffuse regions. For gamma rays, the proximity of the source improves the signal-to-noise ratio, while energy and angular selection suppress astrophysical backgrounds. For neutrinos, directional constraints and reduced atmospheric backgrounds at higher energies aid in isolating the signal. Following we discuss some possible limitations on our current results.
In this study, we adopted a simplified model using a zero-temperature approximation irrespective of the NS’s core temperature. This is a general set-up applied by many literatures to provide bound on DM properties from NS Leane and Tong (2024); Acevedo et al. (2024); Leane et al. (2021); Bose et al. (2022). While higher core temperatures are likely, especially in younger stars, this approach allows us to focus on the primary objectives without introducing the complexities of temperature-dependent factors. Although incorporating the actual core temperature could yield more precise results, it would require additional assumptions and models beyond the scope of our work. We acknowledge that our results may vary by a few orders with improved temperature corrections imposed on NSs leaving room for more detailed future work. Furthermore, in this study, we have opted to use a simplified model where the key parameters, such as the mass of NS, escape velocity, and nucleon form factors, are treated as constants within the NS interior. This approach is intended to provide a manageable framework that offers clearer insights into the core dynamics of our model, allowing for direct comparisons against previous studies for other celestial objects. However, we acknowledge that these quantities vary significantly across the NS interior due to gravitational effects and changes in density, which can influence the capture rate as noted in refs. Bell et al. (2020, 2021d); Anzuini et al. (2021). We also want to point out that, for this work, we use the standardized values for $M_{\rm NS}$ and $R_{\rm NS}$, but the capture rates directly depend on these two parameters. For neutron star mass $M_{\rm NS}$ between 1.4 $M_{\odot}$-2.5 $M_{\odot}$ and $R_{\rm NS}$ from 10-25 km, our reported bounds could vary by nearly one order of magnitude. We are aware of the limitations of this method and plan to address these complexities in future work to more accurately reflect the true nature of NS environments.
With the present formalism, we find a significant bound on the DM scattering cross section from the local NS population. Improved constraints on the $\sigma_{\chi_{n}}$ can be achieved with precise corrections to the NS properties and the nature of DM interactions. The limits on $\sigma_{\chi n}$ obtained for DM mass $m_{\chi}$ are valid for DM with specific scalar-scalar type interactions. Including different DM interaction types will change the relevant bounds in the present work and a detailed study is required. One can also consider the case where DM undergoes direct annihilation instead of annihilation via a long-lived mediator considered in the present work. However, this involves different dynamics of final state annihilation with avenues that can be addressed in future works.
In addition, we want to point out that the precision of our findings strongly hinges on the systematic uncertainties associated with each instrument. Regarding gamma-ray detectors, the Fermi-LAT reveals flux uncertainties of approximately 10%-15%, alongside variations in energy scale (up to 5%-10%) and effective area (around 10%) Ackermann et al. (2012); Abdo et al. (2009). CTA typically experiences flux uncertainties spanning 10%-20%, accompanied by calibration uncertainties of 5%-10% and atmospheric influences ranging from 10%-15% Acharya et al. (2018); Bernlohr (2008). LHAASO estimates flux uncertainties at 15%-20%, with calibration around 10% and atmospheric modeling impacting results by 10%-15% Di Sciascio (2016); Wang et al. (2023). SWGO’s flux uncertainties are approximately 15%-20%, with calibration and atmospheric effects each contributing about 10%-15% Albert et al. (2019). Concerning neutrino detectors, IceCube displays flux uncertainties ranging from 10%-20%, with optical properties of ice introducing uncertainties of up to 10%-15% and calibration uncertainties around 10% Williams (2020); Aartsen et al. (2017c). IceCube Gen 2 demonstrates similar uncertainties, slightly improved to around 10%-15% Aartsen et al. (2021, 2014). KM3Net reports flux uncertainties of 15%-20%, with uncertainties in water properties and calibration reaching up to 10%-15% each Adrian-Martinez et al. (2016); Le Breton (2019). TRIDENT anticipates flux uncertainties of around 15%-20%, with calibration and environmental conditions contributing approximately 10%-15% each Ye et al. (2023); Li et al. (2023). In summary, systematic uncertainties in our chosen detectors may influence our results by approximately 10%-20%.
@@ VII Discussion
In conclusion, this study reveals that present and upcoming neutrino and gamma-ray observatories are not very sensitive to DM-nucleon scattering when a few known local NSs within 0.6 kpc in the neighborhood of Earth are considered and can only provide a weak bound for DM mass $\sim 1$ GeV. Improved limits on DM-nucleon scattering cross section can be achieved with the consideration of about $10^{5}$ local NS population anticipated from the CGW search for a range between low ($\sim$ sub-GeV) to heavy ($\sim$ PeV) DM mass. The corresponding limit on $\sigma_{\chi n}$ with the local NS population is commendable concerning those obtained from DM direct search and astrophysical limits from the Sun and convincing as we compare with indirect bounds from the Galactic NS population. Future CGW searches with better sensitivity are expected to limit further the local NS population, which can further constrain the DM-nucleon scattering with better precision.
@@ Acknowledgements.
A. D. B. acknowledges financial support from DST, India, under Grant No. IFA20-PH250 (INSPIRE Faculty Award). Some part of P. B.’s work was supported by the EOSC Future project which was cofunded by the European Union Horizon Program call INFRAEOSC-03-2020, Grant Agreement No. 101017536. P. B. further acknowledges support from the COFUND action of Horizon Europe’s Marie Sklodowska-Curie Actions research program, Grant Agreement 101081355 (SMASH). The authors would like to thank Dr. Tarak Maity, Dr. Nicholas L. Rodd, and Dr. Sandra Robles for their valuable suggestions.